Evaluate the difference quotient for the given function. Simplify your answer.

f(x) = x2 + 3,    

f(4 + h) − f(4)
h
Answer #1

solution Difference quotient is given f(a+h)-f(x) here x = 4 & fix) = x²+3 *+14+h) = (4+h)²+3 = 16 + + 8h+3 = 62+8h + 19 & f(